THE CO A-X SYSTEM FOR CONSTRAINING COSMOLOGICAL DRIFT OF THE PROTON-ELECTRON MASS RATIO

\underline{M. L. NIU}, E. J. SALUMBIDES, D. ZHAO, J. BAGDONAITE, Department of Physics and Astronomy, and LaserLaB, VU University, De Boelelaan 1081, 1081 HV Amsterdam, The Netherlands; N. DE OLIVEIRA, D. JOYEUX, L. NAHON, Synchrotron Soleil, Orme des Merisiers, St Aubin BP 48, 91192, GIF sur Yvette cedex, France; R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, Massachusetts 02139, USA; and W. UBACHS, Department of Physics and Astronomy, and LaserLaB, VU University, De Boelelaan 1081, 1081 HV Amsterdam, The Netherlands.

The A^{1}Π−X^{1}Σ^{+} band system of carbon monoxide, which has been detected in six highly redshifted galaxies (z = 1.6−2.7), is identified as a novel probe method to search for possible variations of the proton-electron mass ratio (μ) on cosmological time scales. Laboratory wavelengths of the spectral lines of the A-X (v,0) bands for v = 0 − 9 have been determined at an accuracy of Δλ/λ = 1.5 × 10^{−7} through VUV Fourier-transform absorption spectroscopy, providing a comprehensive and accurate zero-redshift data set^{a}. Two-photon Doppler-free laser spectroscopy has been applied for the (0,0) and (1,0) bands, achieving 3 × 10^{−8} accuracy level, verifying the absorption data. Accurate sensitivity coefficients K_{μ} for a varying μ have been calculated for the CO A-X bands, so that an operational method results to search for μ-variation. The data from both experiments were used to perform an improved analysis of the perturbations in the A^{1}Π, v = 0 and 1 levels by vibrational levels in the D^{1}Δ, I^{1}Σ^{−}, e^{3}Σ^{−}, d^{3}Δ, and a^{′3}Σ^{+} states. The object Q1237+064 will be observed in May 2013, aiming for a good signal-to-noise spectrum containing the CO A-X bands as well as the H_{2} Lyman and Werner bands. This should result in an accurate and robust constraint on Δμ/μ.

aE.J. Salumbides, M.L. Niu, J. Bagdonaite, N. de Oliveira, D. Joyeux, L. Nahon and W. Ubachs, Phys. Rev. A 86, 022510